Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: In D configuration (all antennas moved as of the 26th; end of D is 27 December). Missing scan issues associated with changing correlator configs and CBE integration times has not yet been hunted down but can be avoided. 3-bit testing ongoing (report later). EA25 back in the array. 63/64 BlB pairs available. Double maintenance day begins.
  • Key dates:
    • 26 Sep - 27 Dec: D configuration
  • Visitors:
    • Michiel Brentjens: 11 Jun - 11 Nov (Holography; characterizing polarized beams)

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
    • Antennas: EA25 back (used for LSCXKu last night; should be available broadly tonight); EA19 is out.
    • Receivers:
      • EA24 for first L-band solar capable Rx - pushed to 26 Oct.
      • Perley: RF Switch testing ongoing
      • Calibration (Perley): Gain compression/expansion ongoing (Hayward). - work begins again this week.
    • Holography (Brentjens): Notes: Actual campaign started; K band test (with inadvertent excitement).
  • WIDAR (Rupen):
    • Notes (Rupen/Sowinski): Some issues with recent changes to BlBs; removed on Friday. 1 bad BlB pair currently; to be replaced today(?). Hunt to understand the missing scans is ongoing (thus far not able to reproduce reliably). 3-bit sampler tests are highest focus (status):
  • Computing (Robnett): Lustre maintenance outage 1000-1200 MDT, 20 October.

Software status and updates (Butler)

Commissioning Group Reports and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!

Commissioning priorities for the week (McMullin, Chandler); Note: Change for Mondays in D configuration - dedicated to science operations/planned commissioning tests.

  • Mon
    • 1000-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Tue: Maintenance day
    • 0630-1700: Maintenance operations
    • 1800-0900: Science operations and testing
  • Wed: Maintenance day
    • 0630-1700: Maintenance operations
    • 1800-0900: Science operations and testing
  • Thu:
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Fri (not-yet-a-holiday):
    • 0900-0900: Science operations and testing
  • Sat-Sun
    • Science operations + planned testing

Issues Affecting Science Observing:

  • Correlator: [Rupen/Carlson w/ Momjian/Sault] V stokes issues.Status - next week.
  • Correlator: [Pokorny/Rupen/McMullin] time mislabeling in SDM data (non-synchronous operation between correlator dumps and CBE can require additional 1/2 integration time padding for online flags). .Needs close-out testing (Michael-update activation time testing).
  • Correlator: [Clark/Carlson] Delay rocking across bands. (October)
    • Array: [All] Zoo of other effects (some data are very good, some still seeing antenna-based IF/Polarization issue).

Testing This week/soon

  • 3-bit sampler testing
  • Holography
  • RSRO D configuration tests (principally data rate; recirc, solar).
  • OTF test.
  • Flux scale testing (Cyg A)
  • Subarrays: cleanup.
  • Stage 2 mixer testing to relax sub 128 MHz BW grids
  • Correlator: Rupen: S-band shadow testing
  • Baseband edge/filter rolloff.
  • Sub-band dependent fshifts.
  • Array: *; Individual issue sleuthing.

Commissioning Groups

RSRO/ECSO Reports

RSRO A configuration status: https://safe.nrao.edu/wiki/bin/view/EVLA/2011AConfiguration

Individual reports (updates)

Misc:

Demo Science Status

AOB

  • Renaming of the EVLA

-- JosephMcMullin - 2011-10-18
Topic revision: r1 - 2011-10-18, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback